Having trouble can't remember how to work this problem out.
4(5n-7)=32

Answer by issacodegard(60)

Hi, first we distribute the 4 to get,
4*5n-4*7=32
20n-28=32
Add 28 to both sides:
20n-28+28=32+28
20n=60
Divide by 20 on both sides:
20n/20=60/20
n=3